### v3

# Zero-Day Audio DeepFake Detection via Retrieval Augmentation and Profile Matching

Abstract
Modern audio deepfake detectors built on foundation models and large training datasets achieve promising detection performance. However, they struggle with zero-day attacks, where the audio samples are generated by novel synthesis methods that models have not seen from reigning training data. Conventional approaches fine-tune the detector, which can be problematic when prompt response is needed. This paper proposes a training-free retrieval-augmented framework for zero-day audio deepfake detection that leverages knowledge representations and voice profile matching. Within this framework, we propose simple yet effective retrieval and ensemble methods that reach performance comparable to supervised baselines and their fine-tuned counterparts on the DeepFake-Eval-2024 benchmark, without any additional model training. We also conduct ablation on voice profile attributes, and demonstrate the cross-database generalizability of the framework with introducing simple and training-free fusion strategies [^1] This study is supported by JST AIP Acceleration Research (JPMJCR24U3) and the New Energy and Industrial Technology Development Organization (NEDO, JPNP22007). This study was partially carried out using the TSUBAME4.0 supercomputer at the Institute of Science Tokyo..
I Introduction
With the advancement of deep neural networks (DNNs), modern audio deepfake detection (ADD) systems, or so-called spoofing countermeasures (CM), have achieved strong performance in both controlled and complex scenarios. Such progress is demonstrated through established evaluation frameworks like the ASVspoof challenge series [12, 16], as well as recent datasets including SpoofCeleb [10] and DeepFake-Eval-2024 (DE2024) [2]. Meanwhile, recent advances in self-supervised learning (SSL) audio features [3, 8] have notably improved CM detection performance on these datasets, particularly when combined with backend classifiers such as AASIST [9, 15] and simple multi-layer perceptrons [7, 18].
However, current CMs primarily focus on empirical performance. Although they can be updated through fine-tuning, it demands notable computational resources and time, and thus often cause considerable delay in adapting to newly-emerged attacking algorithms, as illustrated in Figure 1. These algorithms are represented by not only research outputs, but also commercial services like ElevenLabs [^2] https://elevenlabs.io. They may not be represented in currently-rolling CM training data during that period. Thus, they are called zero-day attacks and should ideally be addressed promptly without fine-tuning or additional training. This need for training-free ADD methodologies motivates the present study. We address the research question: Can we defend zero-day audio Deepfake attacks promptly without any form of additional fine-tuning?
To answer this question, we propose a simple and comprehensive framework against zero-day attacks, utilizing retrieval augmentation (RA). It utilizes a knowledge database containing reference seen data, and addresses the attacks through knowledge retrieval and ensembling based on k-nearest neighbors (k-NN). The knowledge database stores both representation-level features and score-level predictions from a well-trained SSL-based CM. Additionally, we incorporate speaker and voice-related information extracted from specialized foundation models, and seek whether leveraging such profile matching method can further enhance the detection performance. Both the methods and their related models do not involve any additional re-training or fine-tuning by themselves, enabling prompt reaction and adaptation for robustness to the zero-day attacks. Furthermore, in order to validate and enhance the generalizability of the framework, we introduce AI4T [4] as an alternative evaluation dataset, and extend the framework with two simple, training-free fusion strategies.
Figure 1: Demonstration of the vulnerability of deepfake detection models caused by the time lag between the zero-day attack period and model update. The time progression indicates natural time flow.
Figure 2: Outline of the methodology. The knowledge source data contains input recordings and ground truth binary labels. The lock symbols indicate that both the CM and profile feature extractors remain frozen without fine-tuning. Retrieved vectors can be either CM features, profile features, or both, depending on the strategy detailed in Section II-C. Example number of retrieved utterances is $k=4$ in this outline.
While RA has been applied to ADD, existing methods substantially differ from the approach proposed in this study. [11] maintains a vector data pool and retrieves the vectors of the most closely-matched samples during inference to augment the detection process. [13] employs k-NN to select features for fine-tuning the backend classifier. However, apart from fine-tuning the SSL frontend for ADD, both works involve additional training of backend classifier for feature selection and decision making. This study employs RA by directly using existing CMs that has been trained on massive amount of ADD-related datasets, enabling more broadly applicable solution for advanced SSL-based CMs. For completeness, we also repeat the same proposed method via knowledge database with a model fine-tuned on DE2024 training data, providing supplementary insights and contextual positioning of our approach.
II Knowledge-Based Retrieval Augmentation
II-A Source data: DeepFake-Eval-2024
DE2024 [2] serves as the source dataset for this study, as a representative of real-world deepfakes. It contains 56.5 hours of audio content collected from various social media platforms and deepfake detection services, and includes over 40 languages from more than 80 web sources.
In order to adapt and perform a fair comparison with the existing, widely used CMs, we implement the following adaptation schemes on DE2024: 1) Following the fine-tuning setup in [2], 60% and 40% of the original dataset is used for training and evaluation, respectively. Among the training data, we randomly picked 5% of them as a development set. The ratio of real to fake audios across all splits is approximately 39% vs. 61%, with minor variations observed across different durations; 2) To compare with the established baseline, whose pre-trained SSL-based CM is acquired in this study, we segment both training and evaluation partitions into various lengths ranging from 4 seconds (as reported in [2]) to 50 seconds. The training partition is used to construct the knowledge database, maintaining the same real-to-fake data ratio. The evaluation partition serves as the query data, containing zero-day attacks. They are detailed in next subsection.
II-B Knowledge database construction
As shown in the upper part of Fig. 2, our methodology utilizes a knowledge database to enhance both classification accuracy and feature representation for query recordings that may contain zero-day audio deepfake attacks. Let $\mathcal{K}$ denote a knowledge database of $N$ labeled utterances. For each $i\in\mathcal{K}$, we have the following information extracted and stored.
  • $\mathbf{x}_{i,cm}$: CM feature vector. This vector is extracted from the SSL-based CM system, specifically from the output of the SSL frontend (e.g., Wav2Vec2 [1]). The frame-level features are averaged across all frames to produce a single utterance-level representation vector.
  • $\mathbf{x}_{i,prof}$: Profile feature vector. This vector captures multiple voice and speaker attributes using well-trained neural network models. Various attributes including age, gender, emotion, and voice quality such as pitch and rhythm, are inferred and concatenated to form a comprehensive profile feature vector that describes different aspects of the voice and speaker identity.
  • $y_{i}\in\{0,1\}$: This is ground truth binary label indicating whether the utterance is real (0) or fake (1). These labels are manually annotated and provided in the source metadata.
  • $s_{i}\in(0,1)$: This is the prediction score for knowledge ensemble. The scalar is also obtained from the CM, where the SSL feature vector is passed through the linear layer of the CM, followed by a softmax to produce the score.
The detailed description of the SSL-based CM and the profile feature extraction process is presented in Section III-A.
II-C Retrieval augmentation
After constructing the knowledge database, we proceed to performing RA, shown in the bottom part of Fig. 2. Firstly, at the retrieval stage, given a query utterance $q$, we use the same SSL-based CM and profile extractor as constructing $\mathcal{K}$ to obtain the query CM feature vector $\mathbf{x}_{q,cm}$, query profile feature vector $\mathbf{x}_{q,prof}$, and prediction score $s_{q}$. Note that we do not have the binary ground truth label from the query side.
Denoting the query and knowledge vectors as $\mathbf{x}_{q}$ and $\mathbf{x}_{i}$, the set of the $k$ nearest neighbors $\mathcal{N}_{q}$ is found by selecting the indices of the vectors with the highest similarity scores:
$\mathcal{N}_{q}=\{i_{1},\ldots,i_{k}\}\text{ where }d(\mathbf{x}_{q},\mathbf{x}_{i_{1}})\geq\ldots\geq d(\mathbf{x}_{q},\mathbf{x}_{i_{k}}),$ (1)
and for any index $j\notin\{i_{1},\ldots,i_{k}\}$, we have $d(\mathbf{x}_{q},\mathbf{x}_{i_{k}})\geq d(\mathbf{x}_{q},\mathbf{x}_{j})$. Note that $d(\mathbf{x}_{q},\mathbf{x}_{i})$ is the similarity function. In this study, We use cosine similarity:
[…]
The query and knowledge vectors include respectively either the CM $(\mathbf{x}_{q,cm},\mathbf{x}_{i,cm})$ or the profile feature vectors $(\mathbf{x}_{q,prof},\mathbf{x}_{i,prof})$, depending on the retrieval strategies briefed below:
  • CM-only: Using CM feature vectors solely for retrieval. $\mathbf{x}_{q}=\mathbf{x}_{q,cm}$, $\mathbf{x}_{i}=\mathbf{x}_{i,cm}$.
  • Profile-only: Use profile feature vectors solely for retrieval. Here, $\mathbf{x}_{q}=\mathbf{x}_{q,prof}$, $\mathbf{x}_{i}=\mathbf{x}_{i,prof}$.
  • Hybrid: Retrieving $k_{1}=\lfloor k/2\rfloor$ neighbors using CM features and $k_{2}=\lceil k/2\rceil$ neighbors using profile features. Defining $\text{argtop}_{k}d(.)$ as the set of $k$ values that holds highest values of $d(.)$, we have:
    $\mathcal{N}_{q}^{(cm,k_{1})}$ $=\text{argtop}_{k_{1}}d(\mathbf{x}_{q,cm},\mathbf{x}_{i,cm})$ (3)
    $\mathcal{N}_{q}^{(prof,k_{2})}$ $=\text{argtop}_{k_{2}}d(\mathbf{x}_{q,prof},\mathbf{x}_{i,prof})$ (4)
    $\mathcal{N}_{q}^{(hybrid)}$ $=\mathcal{N}_{q}^{(cm,k_{1})}\cup\mathcal{N}_{q}^{(prof,k_{2})}$ (5)
Note that $|\mathcal{N}_{q}^{(hybrid)}|\leq k$ due to potential overlapping retrieval results between the two involved strategies.
We then proceed to the ensemble stage, fetching the neighbor labels, scores, and feature vectors from $\mathcal{K}$, denoting them as $Y_{q}=\{y_{i_{j}}\}_{j=1}^{k}$, $S_{q}=\{s_{i_{j}}\}_{j=1}^{k}$, and $X_{q}=\{\mathbf{x}_{i_{j}}\}_{j=1}^{k}$ respectively. We employ three simple ensemble strategies to aggregate the retrieved information into a final prediction:
  • Majority Voting (MV): The ensemble score is simply computed by identifying the majority label among the retrieved utterances: $\hat{s}_{q}=\arg\max_{c\in\{0,1\}}\sum_{j=1}^{|\mathcal{N}_{q}|}\mathbb{I}[y_{i_{j}}=c]$, where $\mathbb{I}[\cdot]$ is the indicator function and $c$ is the binary class label. Thus, the prediction score is numerically binary: $\hat{s}_{q}\in\{0,1\}$.
  • Ratio-based Scoring (Ratio): The ensemble score is computed as the proportion of positive samples out of those retrieved: $\hat{s}_{q}=\frac{1}{|\mathcal{N}_{q}|}\sum_{j=1}^{|\mathcal{N}_{q}|}y_{i_{j}}=\frac{|\{j:y_{i_{j}}=1\}|}{|\mathcal{N}_{q}|}$.
  • Score-level Averaging (Avg): The ensemble score is simply computed as the mean average of the retrieved prediction scores: $\hat{s}_{q}=\frac{1}{|\mathcal{N}_{q}|}\sum_{j=1}^{|\mathcal{N}_{q}|}s_{i_{j}}$.
The ensembled score is the output of RA. We try different combinations of retrieval and ensemble strategies as specified in Table I, resulting in $3\times 3=9$ systems (S01 to S09).
4 seconds 10 seconds 13 seconds 30 seconds 50 seconds
[…]
System $\mathcal{N}_{q}$ $\hat{s}_{q}$ $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%) $k$ (%) (%)
S01 CM-only MV 20 15.47 86.46 20 13.96 87.81 10 13.15 88.65 20 13.76 88.00 200 16.96 85.98
S02 Ratio 20 14.62 86.43 100 14.09 86.70 20 12.98 87.96 20 11.21 88.92 20 11.59 87.49
S03 Avg 5 27.27 75.73 5 21.91 78.90 200 20.25 79.16 200 16.56 81.21 100 15.35 82.41
S04 Profile-only MV 100 26.07 78.26 50 24.84 79.15 20 23.02 80.33 20 23.86 79.26 10 23.40 80.34
S05 Ratio 100 25.35 78.22 5 23.59 78.37 5 22.18 79.16 5 22.17 79.52 10 22.75 78.55
S06 Avg 50 35.36 68.80 200 31.21 69.13 20 28.06 71.07 5 25.39 75.08 5 30.42 76.39
S07 Hybrid MV 200 15.70 86.38 100 14.69 88.01 100 14.63 87.96 20 13.08 89.07 20 13.87 88.52
S08 Ratio 20 14.78 86.84 10 12.81 87.73 10 12.81 87.52 20 11.25 89.38 20 10.92 88.90
S09 Avg 50 33.19 69.20 200 31.37 69.35 50 27.93 71.89 100 31.19 70.17 200 32.77 69.61
B00 - - - 27.76 75.15 - 22.21 78.95 - 20.36 79.14 - 16.67 81.00 - 14.92 82.22
B01 - - - 12.52 87.69 - 11.11 90.95 - 10.21 90.68 - 9.53 90.91 - 9.96 90.53
TABLE I: Performance of the proposed RA methods on DE2024 evaluation data. $k$ denotes the optimal number of retrieved utterances. Numbers in bold and underline indicate the best and second-best results under each condition, respectively.
III Experimental Setup
III-A Self-supervised learning models
Countermeasures. We use an open-sourced SSL-based CM model [^3] https://huggingface.co/nii-yamagishilab/xls-r-2b-anti-deepfake for this study. The model uses a Wav2Vec2.0 frontend [1] to extract feature vectors, followed by a backend classifier with adaptive pooling, a fully-connected MLP (without bias), and softmax operation for prediction scores. The model was post-trained [^4] As described in [7], post-training is an intermediate supervised training phase for SSL models, with specifically collected data and training schemes to adapt them for certain task. Here, the task is ADD. on approximately 74,000 hours of real and fake audio clips from various datasets. Further details of model architecture and training are described in [7].
Profile extractor. We employ vox_profile [^5] https://github.com/tiantiaf0627/vox-profile-release [6] to extract profile feature vectors using well-trained foundation models. We construct the features by concatenating four attributes: 1) Age: a scalar representing age value in terms of 10 uniformed intervals from 0 to 100 years old666e.g., A value of 7 means an estimated age of 61-70 years old.; 2) Gender: a 2-dimensional one-hot vector for male/female classification; 3) Emotion: a scalar indicating the emotion trait (anger, fear, neutral, etc.) and 256-dimensional embedding vector extracted from the intermediate of the foundational model; 4) Voice quality: a 25-dimensional vector capturing five perceptual aspects: pitch, volume, clarity, rhythm, and voice texture, following the taxonomy defined in [5]. More details about how the representations are extracted can be found in [6].
III-B Implementation & evaluation
Baselines. To demonstrate the effectiveness of RA against fine-tuning, we compare it with two baselines: 1) The post-trained model that returns $s_{q}$ from the DE2024 evaluation data; 2) A fine-tuned model from the post-trained one using the DE2024 training data (same data used for knowledge database of RA). For fine-tuning, we use a learning rate of $1e^{-7}$ and other settings consistent with [7]. We fine-tune the post-trained model for 10 epochs and infer $s_{q}$ for evaluation, with notable improvements from the baseline, as reported in next section. The baseline and fine-tuned models are denoted as B00 and B01, respectively in the rest of this paper.
Hyperparameters. The only hyperparameter is the number of utterances $k$ to be retrieved from the knowledge database. We range $k$ from 5 to 200 at several discrete values. For each system, we report results using the optimal $k$ value based on performance on the development set.
Hardware. All SSL model inference and fine-tuning experiments are performed on a computing node with eight AMD EPYC 9654 CPUs and one NVIDIA H100 GPU. Knowledge retrieval, ensemble, and evaluation use the same node without involving the GPU resources.
Evaluation data. For validating the proposed RA framework itself, we use the DE2024 evaluation set under the corresponding various duration conditions. For the extension covered in Section V, we additionally acquire AI4T [4], a social-media audio Deepfake corpus with 13 hours of speech in 8 languages. We follow the official protocol described in [4] [^7] https://github.com/davidcombei/AI4T/tree/main to construct the evaluation data. This dataset differs substantially from DE2024 in its attacking methods, which makes it suitable for benchmarking generalizability.
Evaluation metrics. We acquire two metrics: 1) Equal error rate (EER), following standard practice in ADD; 2) Accuracy with a fixed threshold of 0.5. A fixed threshold is used instead of one optimized on a development set because zero-day attacks may involve new, unseen artifacts, making it difficult to derive reliable thresholds from existing data, similar to the limitations of model fine-tuning in real-world scenarios.
IV Results and Analysis
IV-A Main Results
Results for different RA strategies are presented in Table I. There are several key findings which are listed below.
Retrieval augmentation methods which directly incorporate class labels for knowledge ensemble are effective to zero-day attacks, being comparable to fine-tuned models. The proposed methods with MV and ratio-based scoring ensemble consistently outperform B00 in terms of accuracy. Among them, the ratio-based ensemble with a hybrid retrieval strategy (S08) achieves performance comparable to B01, having either the 2nd best EER or accuracy (or both) across all duration conditions. This indicates the effectiveness of the proposed information acquisition mechanism. In terms of EER, while the best results are generally achieved by the fine-tuned model, the ratio-based fusion using either the CM feature alone or both features (S02, S08) achieves nearly comparable performance, with the gap further narrowing for longer durations. Meanwhile, score averaging often fails to surpass the baseline. The above findings demonstrate the efficacy of RA for ADD as a binary classification task.
The profile information does not work well independently, but is sufficient compensation for the CM features. Regarding the retrieval method, the analysis and results show that using CM-only or Hybrid for retrieval, with Ratio as ensemble method (S02, S08), achieved comparable performance to that of the fine-tuned models. S08 reaches relatively more stably comparable performance against B01 across all duration conditions, which incorporates the profile information. In contrast, using only profile features (S04-S06) do not return as promising performance. This is likely because the foundation models used for profile feature extraction mainly capture voice and speaker characteristics rather than real/fake cues, as discussed in Section III-A. Nevertheless, the strong performance of S08 suggests that speaker- and voice-related information can effectively complement CM-based representations. Section IV-B examines the individual contribution of these components.
Apart from the above main findings, it may be also worth noting that audio duration may have an impact on the relative efficacy of RA. The relative performance gap between (especially) B00, the best of the proposed methods, and B01 decrease with longer audio segments. This may be due to more informative embeddings, making models less sensitive to different strategies. Investigating these duration-dependent effects is a possible direction for future work.
4 seconds 30 seconds
[…]
S08 14.78 86.84 11.25 89.38
w/o Age & Gender 14.53 87.06 12.06 88.76
w/o Emotion 14.48 86.88 12.15 88.76
w/o Voice Quality 16.14 85.54 13.07 87.05
TABLE II: Probing attributes from profile feature extractor (w/o: without). Number of retrieved utterances here is $k=20$.
IV-B Ablation on voice profile attributes
Earlier results shown in last section demonstrates that the profile features are good compensation. It is then natural to additionally analyze, among the several attributes involved in constructing the profile features, which ones drive to positive results more than others. We use S08 to compare the attributes by probing them in the profile features: age & gender, emotion, and voice quality. We remove each attribute individually to assess its contribution to performance.
Results are presented in Table II. While properties more closely related to speakers, such as age, sex and emotion, have a relatively weaker impact, removing voice quality more substantially deteriorates the performance in both metrics, especially for 4-second condition. This indicates the relatively higher importance of voice qualities, which can be affected by not only the speaker, but also the surrounding environment and context, considering the nature of the DE2024.
V Cross-database evaluation
In the previous sections, the knowledge retrieval source, fine-tuning data, and evaluation sets were all derived from DE2024. Despite its in-the-wild nature, this “in-domain” setting does not fully show how well the proposed RA generalize to other databases.
Therefore, in this section, we keep the knowledge source for RA (and fine-tuning data) fixed to DE2024 (4-second condition), and additionally include AI4T as part of evaluation, as mentioned in Section III-B. We further extend the RA framework by testing two simple fusion methods that may improve cross-database generalizability. We also evaluate several previously proposed countermeasures [9, 17], together with their fine-tuned variants, where fine-tuning is performed on the DE2024 training data.
V-A Linear and Selective fusion
From the results shown in Table I, it was found that the proposed method (e.g. S08) is training-free and holds promising performance, while the fine-tuned system still performs moderately better. These two approaches may provide complementary strengths.
Therefore, we evaluate two simple fusion strategies here: 1) Linear fusion, where we consider simple fusion between $s_{q}$ and $\hat{s}_{q}$ shown in Section II-C by $s_{linear}=\lambda*s_{q}+(1-\lambda)*\hat{s}_{q}$; 2) Selective fusion, where the query representation is firstly input to a specialized module for deciding whether it is in-domain or out of domain (OOD), then routed to the most suitable detector. If the query is judged as in-domain, we directly use the score from the SSL-based CM. Conversely, if the query is judged to be OOD, we switch to the proposed RA method with its knowledge database, and use the ensembled score as the final output.
We employ a training-free $k$-NN based OOD that measures cosine similarity between input embedding and one in a knowledge database [14]. For better understanding, the framework of selective fusion is illustrated in Fig. 3. Here, we utilize two knowledge databases with distinct roles: 1) A knowledge database dedicated to the training-free OOD detection; 2) Another database used for RA of the proposed deepfake detector.
Figure 3: Selective fusion of the proposed training-free deepfake detector and SSL-based CM. The lock symbols indicate the models that are remain frozen without fine-tuning. The text in italian refers to the source data to construct the knowledge databases. We use B00 and S08 as example systems here.
DE2024 (4 seconds) AI4T
[…]
No fine-tuning with DE2024 training data
AASIST [9] - 55.73 37.71 - 55.81 57.30
P3 [17] - 47.83 56.15 - 53.45 56.26
B00 - 27.76 75.15 - 10.48 75.51
S08 - 14.78 86.84 - 16.98 83.24
B00 ++ S08 - 13.78 87.20 - 15.92 84.06
B00 $\lor$ S08 0.76 14.22 87.16 1.00 16.98 83.24
Fine-tuned with with DE2024 training data
AASIST - 18.87 82.50 - 53.91 44.56
P3 - 14.04 86.07 - 17.91 67.92
B01 - 12.52 87.69 - 8.94 90.20
S08* - 12.56 89.11 - 14.59 85.64
B01 ++ S08* - 12.02 89.65 - 13.29 87.12
B01 $\lor$ S08* 0.62 12.25 89.19 0.43 11.71 89.02
TABLE III: Results on the cross-database evaluation. $\mathcal{R}$ denotes the proportion of OOD data classified from the query dataset (1.00 means all data are OOD). Numbers in bold indicate the best results.
V-B Implementation details
Data. Since they include both outputs from SSL-based CM and RA, we consider two system pairs: 1) B00 and S08, which are based on the SSL-based CM without fine-tuning; 2) B01 and S08*, which are from the SSL CM fine-tuned on DE2024 training data (S08* represents RA built on B01). In terms of the source data of the two knowledge databases, the knowledge database for the OOD detection is constructed from a gender- and language-balanced subsample of the original training data used to post-train the SSL-based CM [^8] Details of the training data, including statistics and attribute-wise distributions, can be found in https://github.com/nii-yamagishilab/AntiDeepfake.. The size of the database is 26,000 samples, which is almost same as the DE2024 4-second training data. To set the OOD decision threshold, we further construct a small development set by randomly sampling 1,000 utterances from either knowledge database. Among the RA configurations, we use S08 with the knowledge database described in Section II-C.
Implementation. For linear fusion, we need to determine the weight value $\lambda$ between the SSL CM scores and the RA ensembled scores on DE2024 development set, and use such $\lambda$ for all query datasets involved. In this section, $\lambda=0.1$ for both fusion methods. For the $k$-NN based OOD detection, $k$ was set to 20. The linear and selective fusion methods of two systems $A$ and $B$ are correspondingly denoted with symbols $A+B$ and $A\lor B$ in Table III, respectively.
V-C Results
The results are summarized in Table III. We analyze the results separately depending on whether the backbone SSL-based CM model is fine-tuned with the DE2024 training data or not.
No fine-tuning with DE2024 training data. On the DE2024 evaluation data, S08 substantially improves over the baseline B00 and outperforms prior works [9, 17]. The linear fusion B00++S08 resulted in the best EER and accuracy. However, on the more-mismatched AI4T, S08 exhibits rather not promising zero-shot generalization in terms of EER compared to the direct application of the SSL-based CM. This indicates that the proposed RA-based deepfake detection method (S08) is training-free, but suffers from the domain mismatch between the knowledge database based on the DE2024 training data and AI4T. The linear fusion improves the accuracy over using S08 alone, but yields worse EER.
Fine-tuning with DE2024 training data. Even if the models are fine-tuned with the DE2024 training data, the general tendency is the same as the above non fine-tuning case. On the DE2024 evaluation, S08 achieves performance comparable to the fine-tuned SSL-based CM B01 and outperforms prior works [9, 17]. The linear fusion B01 ++ S08* achieves the best overall performance on DE2024, reaching the lowest EER (12.02%) and highest accuracy (89.65%). Meanwhile, on AI4T, S08 is not the good choice anymore and the baseline B01 resulted in the best performance in terms of both the accuracy and EER. These observations indicate that even training-free approaches leveraging knowledge databases may be affected by domain mismatch, and addressing this will likely be an interesting and important research direction for future work.
VI Conclusion
In this study, we have addressed zero-day attacks in audio deepfake detection, where systems must use existing knowledge to rapidly handle unseen attacks. We have introduced an retrieval augmentation framework based on knowledge databases and various related strategies, without any manner of fine-tuning or additional training. The framework uses both state-of-the-art CM and voice profile representations. Leveraging DE2024 with various duration conditions, we have found that the proposed method, especially when combining SSL and profile features from the retrieved data, achieves competitive performance.
Finally, we have evaluated the framework on the cross-database scenario and show that the even training-free approach leveraging knowledge databases may be affected by domain mismatch. Future work may explore acquiring temporal information via frame-level SSL features, modeling profile information more effectively, and extending the framework to multi-classification ADD-related tasks.